2-114. Determine the angle 0 between the two cords 3 m B. 4 m 6 m 2. m -3 m Prob. 2-114 C.
Answer #1

LETALS 2-114. 3 m 3m OST- t y 3mM In DAOB, AB = 1 OA²+ OB² = 3 +42 = 5m AB= 5 m In DAMN, AM = JAN*Max = J 3²+32 = STB In AACM